OpenStudy (anonymous):

4th root of: (10e^(iπ/2)) CALCULATE and GIVE YOUR ANSWER in CARTESIAN FORM!

OpenStudy (dumbcow):

\[(10e^{i \pi/2})^{1/4} = \sqrt[4]{10} e^{i \pi/8} = \sqrt[4]{10}(\cos \frac{\pi}{8} +i \sin \frac{\pi}{8})\]

OpenStudy (dumbcow):

do you need all 4 roots?

OpenStudy (anonymous):

well i need a final answer

OpenStudy (anonymous):

so i guess yes. because i need any info to write in Cartesian form

OpenStudy (dumbcow):

cartesian form is just a+bi a = rcos b = rsin

OpenStudy (anonymous):

ok so then how would i write my final answer

OpenStudy (dumbcow):

\[r(\cos \theta +i \sin \theta) ^{1/4} = r^{1/4} (\cos \frac{\theta +2\pi k}{4} +i \sin \frac{\theta +2\pi k}{4})\] k = 0,1,2,3 this gives 4 solutions in cartesian form: \[\sqrt[4]{10} \cos \frac{\pi/2 +2\pi k}{4} +\sqrt[4]{10}\sin \frac{\pi/2 +2\pi k}{4} i\]

OpenStudy (anonymous):

thanks for your help

OpenStudy (dumbcow):

your welcome here is calculator to verify http://www.wolframalpha.com/input/?i=%2810e%5E%28i+pi%2F2%29%29%5E1%2F4